# The Standard Model lifetime is slightly shorter

### v1

@@ Appendix B Transverse gauge modes
Let us evaluate the action of the $L_{Ai}+L_{Bi}=L_{Ci}$ operator on spherical harmonics in (25). Because $A$ and $B$ sectors commute, the $C$ operators themselves form an $SU(2)_{C}$ algebra that represents a simultaneous rotation in both $SU(2)_{A,B}$ spaces. Its states can be labeled by $l_{C}$ that ranges from $|l_{A}-l_{B}|\leq l_{C}\leq l_{A}+l_{B}$. The hyperspherical harmonics $Y_{jm_{A}m_{B}}$ correspond to $(j/2,j/2)=(l_{A},l_{B})$. Because $l_{A}=l_{B}$, the minimal value of $l_{C}$ is always 00 for each fixed $j$ in (25). The eigenvalue of the Casimir operator $L_{C}^{2}=\sum_{i}\left(L_{Ai}+L_{Bi}\right)^{2}$ is in general given by $l_{C}(l_{C}+1)$ (as for the canonical $SU(2)_{C}$) and therefore vanishes for $l_{C}=0$. This means that there will always be a single zero eigenvalue in the Gram matrix of $T^{1}$, the rest are non-zero. Without the zero, the matrix would be full rank with the degeneracy of $(j+1)^{2}$, because none of the states would be lost in the last equality of (25). However, the zero is present for each $j$ (this is in contrast to (22), where the zero is there only for $j=0$) and therefore reduces $d_{j}^{(T1)}$ by 1 and we have
@@
Additional overlaps among $T^{1,2}$ transverse modes. We want to prove here that the transverse modes $T^{1,2}$, defined in (20), do not span the whole $2j(j+2)$ dimensional space of transverse fluctuations. Specifically, at each $j$ we find that precisely $j$ modes are linearly dependent, such that
@@
Firstly, we are going to prove that all $T^{1}$ and $T^{2}$ modes belonging to the first class (with $m_{A}+m_{B}\neq 0$) are linearly independent. With a little bit of algebra it can be shown that the $T^{1,2}$ modes take on the following form
@@
$i\sum_{\sigma=-\frac{j}{2}}^{m}T^{2;j\sigma(-\sigma)}_{\mu}=\begin{pmatrix}iP_{m}\\ Q_{m}\end{pmatrix}$ (34)
@@
Appendix C Canonical basis of transverse modes
@@
Let us comment on the mechanics behind (38). The main observation is that, under a rotation acting on the vector field $T^{\pm}_{\mu}(\hat{x})$, it can be shown that the combination $\epsilon_{\mu\nu\rho\sigma}(\tilde{\sigma}_{\nu})_{m_{s_{A}}m_{s_{B}}}L_{\rho\sigma}Y_{jm_{l_{A}}m_{l_{B}}}(\hat{x})$ transforms as the tensor product $(j/2,j/2)\otimes(1/2,1/2)$, where the first and second factor act respectively on the indices $m_{l_{A}},m_{l_{B}}$ and $m_{s_{A}},m_{s_{B}}$. In order to project on the $((j\pm 1)/2,(j\mp 1)/2)$ subrepresentations, the ones characterizing transverse modes, we contract with $C^{\frac{j\pm 1}{2}}C^{\frac{j\mp 1}{2}}$. This is completely analogous to the standard procedure with the rotation group $SO(3)$, with the difference here that there are two factors of $SO(3)\simeq SU(2)$, so we need to multiply by two Clebsch-Gordan coefficients. As a cross check of the consistency of the method, it can be verified that contracting with $C^{\frac{j+1}{2}}C^{\frac{j+1}{2}}$ or with $C^{\frac{j-1}{2}}C^{\frac{j-1}{2}}$ gives zero, because diagonal modes are already projected out by the $\epsilon_{\mu\nu\rho\sigma}$. The modes of (38) are proportional to the ${\cal T}^{\pm}$ modes presented in the main text.
Appendix D Counting in orbital modes
@@
We saw that the transverse modes in either basis, ${\cal T}_{j}^{\pm}$ or $T^{i=1,2,3}_{\mu}$, have $l=j$. Their degeneracy is $2l(l+2)=2j(j+2)$. The other two modes with $l=j$ are

### v2

@@ II Transverse gauge modes
Let us evaluate the action of the $L_{Ai}+L_{Bi}=L_{Ci}$ operator on spherical harmonics in (30). Because $A$ and $B$ sectors commute, the $C$ operators themselves form an $SU(2)_{C}$ algebra that represents a simultaneous rotation in both $SU(2)_{A,B}$ spaces. Its states can be labeled by $l_{C}$ that ranges from $|l_{A}-l_{B}|\leq l_{C}\leq l_{A}+l_{B}$. The hyperspherical harmonics $Y_{jm_{A}m_{B}}$ correspond to $(l_{A},l_{B})=(j/2,j/2)$. Because $l_{A}=l_{B}$, the minimal value of $l_{C}$ is always 00 for each fixed $j$ in (30). The eigenvalue of the Casimir operator $L_{C}^{2}=\sum_{i}\left(L_{Ai}+L_{Bi}\right)^{2}$ is in general given by $l_{C}(l_{C}+1)$ (as for the canonical $SU(2)_{C}$) and therefore vanishes for $l_{C}=0$. This means that there will always be a single zero eigenvalue in the Gram matrix of $T^{1}$, the rest are non-zero. Without the zero, the matrix would be full rank with the degeneracy of $(j+1)^{2}$, because none of the states would be lost in the last equality of (30). However, the zero is present for each $j$ (this is in contrast to (27), where the zero is there only for $j=0$) and therefore reduces $d_{j}^{(T1)}$ by 1 and we have
@@
Additional overlaps among $T^{1,2}$ transverse modes. We want to prove here that the transverse modes $T^{1;j}$ and $T^{2;j}$, defined in (25), do not span the whole $2j(j+2)$ dimensional space of transverse fluctuations. Specifically, at each $j$ we find that precisely $j$ modes are linearly dependent, such that
@@
Firstly, we are going to prove that all $T^{1}$ and $T^{2}$ modes belonging to the first class with $m_{A}+m_{B}\neq 0$ are linearly independent. With a little bit of algebra it can be shown that the $T^{1,2}$ modes take on the following form
@@
$i\sum_{\sigma=-\frac{j}{2}}^{m}T^{2;j\sigma(-\sigma)}_{\mu}=\begin{pmatrix}iP_{m}\\ Q_{m}\end{pmatrix}\,,$ (39)
@@
III Canonical basis of transverse modes
@@
Let us comment on the mechanics behind (43). The main observation is that, under a rotation acting on the vector field $T^{\pm;j}_{\mu}(\hat{x})$, it can be shown that the combination $\epsilon_{\mu\nu\rho\sigma}(\tilde{\sigma}_{\nu})_{m_{s_{A}}m_{s_{B}}}L_{\rho\sigma}Y_{jm_{l_{A}}m_{l_{B}}}(\hat{x})$ transforms as the tensor product $(j/2,j/2)\otimes(1/2,1/2)$, where the first and second factor act respectively on the indices $m_{l_{A}},m_{l_{B}}$ and $m_{s_{A}},m_{s_{B}}$. In order to project on the $((j\pm 1)/2,(j\mp 1)/2)$ subrepresentations, the ones characterizing transverse modes, we contract with $C^{\frac{j\pm 1}{2}}C^{\frac{j\mp 1}{2}}$. This is completely analogous to the standard procedure with the rotation group $SO(3)$, with the difference here that there are two factors of $SO(3)\simeq SU(2)$, so we need to multiply by two Clebsch-Gordan coefficients. As a cross check of the consistency of the method, it can be verified that contracting with $C^{\frac{j+1}{2}}C^{\frac{j+1}{2}}$ or with $C^{\frac{j-1}{2}}C^{\frac{j-1}{2}}$ gives zero, because diagonal modes are already projected out by the $\epsilon_{\mu\nu\rho\sigma}$. The modes of (43) are proportional to the ${\cal T}_{j}^{\pm}$ modes presented in the main text.
IV Counting in orbital modes
@@
We saw that the transverse modes in either basis, ${\cal T}_{j}^{\pm}$ or $T^{i;j}_{\mu}$, have $l=j$. Their degeneracy is $2l(l+2)=2j(j+2)$. The other two modes with $l=j$ are